What is the molar mass of copper (lll) nitrate (Cu(NO3)3
1
Expert's answer
2019-06-05T05:02:45-0400
"Mollar mass( Cu(NO_3)_3)= 65.546 +14.0067\times1\times3 + 15.9994\times 3\times 3 = 249.5607 = 249.56 \frac{g}{mol}"